OpenStudy (anonymous):

Use elimination to solve. 2x-3y=1 5x+4y=14

OpenStudy (anonymous):

Well, first you would have to figure out which values x and y can both have but still be equal. Do you have any in mind yet?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

2 to take the place of x, or y?

OpenStudy (anonymous):

X?

OpenStudy (anonymous):

Yes, if 2 was to take the place of x, it would make sense and be equal to the answer put. If you used 2 for x, It would look like this: (2*2) - 3y = 1 (5*2) + 4y = 14

OpenStudy (anonymous):

So that would mean that y would have to be equal to 1.

OpenStudy (anonymous):

Do you understand now? :)

OpenStudy (anonymous):

2x - 3y = 1 5x + 4y = 14 2x2 - 3 = 1 2x5 + 4 = 14

OpenStudy (anonymous):

How are they equal though ? It's 4-3y=1 and 10+4y=14

OpenStudy (anonymous):

You are trying to find the values of x and y, correct?

OpenStudy (anonymous):

Do x and y both have to be the same in your math problem that was assigned to you?

OpenStudy (anonymous):

No I have to find an x and y that make the equations equal

OpenStudy (anonymous):

You mean to find x and y values so that they equal the same answer?

OpenStudy (anonymous):

Yeah exactly

OpenStudy (anonymous):

Oh, ok. well, what have we found out so far? Do you have an exact answer that they need to end as, or are we just trying to find out one?

OpenStudy (anonymous):

Fund as

OpenStudy (anonymous):

Example, did your teacher, or whomever assigned this to you give you an answer that they both must end in? Or are you supposed to figure it out yourselves?

OpenStudy (anonymous):

Just find an x and y power in each problem that equal 1 and 14

OpenStudy (anonymous):

Ok, well I see that x should probably equal something bigger than 2 to make it a large enough number to be able to equal y.

OpenStudy (anonymous):

Ok, so x and y have to both equal 14 as well as 1? in each problem?

OpenStudy (anonymous):

Yeah like the first problem equals 1so the x and y need to help it equal one and 14 in the 2nd problem

OpenStudy (anonymous):

ok, well, the second problem is 5x + 4 = 14, and we need to find out how to make it equal 1. So, we would have to use a negative number as x then.

OpenStudy (anonymous):

X wouod equal 2

OpenStudy (anonymous):

ok. but how would that equal to 1? y would have to be a negative number. Maybe y could be -2

OpenStudy (anonymous):

That would equal positive 2, and there is no other possible way that we will be able to make it equal 1 unless we use a decimal, and that wouldn't even equal 1 exactly because it would equal somewhere around 1 as a decimal.

OpenStudy (anonymous):

Are you sure it asks you to find how to make x and y values to equal both 1 and 14?

OpenStudy (anonymous):

Do you know hoe to make those drawings one ere? I'll write it out

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

How do you do the writing ?

OpenStudy (anonymous):

Click, the draw button below this box where you type in a 'comment'. There are 3 blue buttons, click where it says 'Draw'.

OpenStudy (anonymous):